OpenStudy (anonymous):

MEDAL AND FAN!!! A man’s suit was on sale for 15% off the regular price. The sale price was $216.75. What was the regular price?

OpenStudy (anonymous):

Here are the choices: A. $249.26 B. $250 C. $225 D. $255

OpenStudy (anonymous):

ok i think this might b C or d but im not sure u have to tag someone else

OpenStudy (tkhunny):

Name the regular price.

OpenStudy (anonymous):

It does not give the regular price.

OpenStudy (anonymous):

Youre supposed to find it

OpenStudy (tkhunny):

That's why I told you to NAME it. Call it something so we can talk about it.

OpenStudy (tkhunny):

Seriously, call it "Steve", if you like. Just give it a name.

OpenStudy (anonymous):

Okay, so this is a percent problem. The reduced price is $216.75 as we know. Are you asking me to choose on of the answer choices as an answer then we talk about it? If you are then I'm going to pick $250.00

OpenStudy (tkhunny):

#1 thing to do. If you don't know it's value, give it a name. Call it something. Once we have a name, we can write equations and create meaningful expressions.

OpenStudy (anonymous):

Okay I think we will go with "Drew".

OpenStudy (tkhunny):

Fine. Drew = Original Purchase Price. What is the discount? The discount percentage is 15%. Write an expression that represents the discount.

OpenStudy (anonymous):

I'm not sure how to do that. D= 216.75

OpenStudy (tkhunny):

How about Drew * 15% = 0.15 * Drew This is the Discount. Do you believe?

OpenStudy (anonymous):

Yes. Okay 15% = 0.15 * Drew.

OpenStudy (tkhunny):

Okay, now we need an expression for the final purchase price. Drew is the Original Purchase Price 0.15 * Drew is the Discount Write an expression that represents the Sale Price.

OpenStudy (anonymous):

D = $216.75 I hope this will suffice.

OpenStudy (tkhunny):

That's no good. That IS the final sale price. We need an expression containing the Original Price that REPRESENTS the final sale price.

OpenStudy (anonymous):

I have no idea how to do that. All I could tell you about what your asking me is that it has Drew in it somewhere.

OpenStudy (tkhunny):

Yes you do. You do it all the time at the store. If the price is $100 and the discount is $10, the sale price is $100 - $10 = $90. Just do that with Drew.

OpenStudy (anonymous):

Okay. I'm going to try it. $Drew - 15% = $216.75

OpenStudy (tkhunny):

Drew - 0.15*Drew = $216.75 That's it!!! Good work. How do we solve?

OpenStudy (tkhunny):

gtg Drew - 0.15*Drew = $216.75 Distributive Property Drew*(1 - 0.15) = $216.75 Resolve Parentheses Drew*(0.85) = $216.75 Divide by 0.85 Drew*(0.85)/0.85 = $216.75/0.85 Drew*(1) = $216.75/0.85 Drew = $216.75/0.85 = $255.00 and we are done. Once we get the name, we are off and running. Without the name, it is quite laborious. I do recommend picking simpler names. I noticed you kept using "D". That is much simpler than "Drew". It doesn't matter much, though. Just pick names you can remember and write a good definition.

OpenStudy (anonymous):

Well thanks.